Two runners, one averaging 8 miles per hour and the other 6 miles per hour, they start at the same place and run along the same trail. The slower runner arrives at the end of the trail a half hour after the faster runner. How far did each person run?
Answer by ankor@dixie-net.com(22740)

Two runners, one averaging 8 miles per hour and the other 6 miles per hour, they start at the same place and run along the same trail.
The slower runner arrives at the end of the trail a half hour after the faster runner.
How far did each person run?
:
Find their times first
:
t = 8mph runner time
(t+.5) = 6mph runner time
:
they both ran the same distance, write a distance equation; dist = speed * time
8t = 6(t+.5)
8t = 6t + 3
8t - 6t = 3
2t = 3
t = 3/2
t = 1.5 hrs
find the dist
8(1.5) = 12 mi
:
check the dist using the 2nd runner time
6(1.5+.5) = 12 mi